OpenStudy (anonymous):

How would I solve this for zero? -x (1-x) = 0

OpenStudy (coconutjj):

Use distributive property

OpenStudy (coconutjj):

solve quadratic equation

OpenStudy (nikato):

if you're solving for x, you can seperate to 2 equations when -x=0 and 1-x=0

OpenStudy (anonymous):

the question is when this equation become equal zero -x(x-1)=0 so this equation equal zero either x=0 -(0)(0-1)=0, -(0)(-1)=0 or x=1 -(1)((1)-1)=0, -(1)(1-1)=0, -1(0)=0 so the answer 0 and 1